\begin{thebibliography}{99}


\bibitem{CS} L. Collatz, U. Sinogowitz, Spektren endlicher grafen, Abh. Math. Sem. Univ. Hamburg. 21(1957) 63--77.


%\bibitem{CCZ} Sarula Chang, An Chang, Yirong Zheng, The leaf-free graphs with nullity $2c(G)-1$,  Discrete Appl. Math. 277(2020) 44--54.

%\bibitem{CT} Li Chen, Fenglei Tian, Skew-rank of an oriented graph with edge-disjoint cycles, Linear and Multilinear Algebra 64(2016) 1197--1206.

\bibitem{CHL} Chen Chen, Jing Huang, Shuchao Li, On the relation between the H-rank of a mixed graph and the matching number of its underlying graph. Linear Multilinear Algebra 66(9)(2018) 1853--1869.

%\bibitem{DMH} D.Cvetkovic, M.Doob, H.Sachs, Spectra of graphs. New York(NY): Academic Press; 1980.

\bibitem{FDD} Yizheng Fan, Wenxue Du, Chunlong Dong, The nullity of bicyclic signed graphs, Linear Multilinear Algebra 62(2014) 242--251.

\bibitem{FYZW} Yizheng Fan, Long Wang, Bounds for the positive and negative inertia index of a graph, Linear Algebra Appl. 522(2017) 15--27.

\bibitem{FWW} Yizheng Fan, Yue Wang, Yi Wang, A note on the nullity of unicyclic signed graphs, Linear Algebra Appl. 438(2013) 1193--1200.

\bibitem{FHLL} Zhimin Feng, Jing Huang, Shuchao Li, Xiaobing Luo, Relationship between the rank and the matching number of a graph, Appl. Math. Comput.  354(2019) 411--421.

%\bibitem{GR} G. Godsil, G. Royle, Algebraic Graph Theory, Graduate Texts in Mathematics, vol. 207, SpringerVerlag, 2001.



\bibitem{HHD} Shengjie He, Rong-Xia Hao, Fengming Dong, The rank of a complex unit gain graph in terms of the matching number, Linear Algebra Appl. 589(2020) 158--185.

 \bibitem{HHL} Shengjie He, Rong-Xia Hao, Hong-Jian Lai, Bounds for the matching number and cyclomatic number of a signed graph in terms of rank, Linear Algebra Appl. 572(2019) 273--291.

\bibitem{HHY} Shengjie He, Rong-Xia Hao, Aimei Yu, Bounds for the rank of a complex unit gain graph in terms of the independence number, Linear Multilinear Algebra, 2020. https://doi.org/10.1080/03081087.2020.1761768.

%\bibitem{HLP} Yaoping Hou, Jiongsheng Li, Yongliang Pan, On the Laplacian eigenvalues of signed graphs, Linear Multilinear Algebra 51(2003) 21--30.

\bibitem{LiW} Shuchao Li, Ting Wang, On the relation between the adjacency rank of a complex unit gain graph and the matching number of its underlying graph, Linear Multilinear Algebra. 2020. https://doi.org/10.1080/03081087.2020.1776205.

\bibitem{LGUO} Xin Li, Ji-Ming Guo, No graph with nullity $\eta(G)=|V(G)|-2m(G)+2c(G)-1$, Discrete Appl. Math. 268(2019) 130--136.

%\bibitem{LXWX} Xueliang Li, Wen Xia, Adjacency rank and independence number of a signed
%graph, Bull. Malays. Math. Sci. Soc. 43(2020) 993--1007

\bibitem{LXYG} Xueliang Li, Guihai Yu,  The skew-rank of oriented graphs, Sci. Sin. Math. 45(2015) 93--104. (in Chinese)

%\bibitem{LH} Yu Liu, Lihua You, Further results on the nullity of signed graphs, J. Appl. Math. 2014(2014) 483735, 8pp.

%\bibitem{LUWH} Lu Lu, Jianfeng Wang, Qiongxiang Huang, Complex unit gain graphs with exactly one positive eigenvalue,  608(2021)   270--281.

\bibitem{LWX} Yong Lu, Ligong Wang, Peng Xiao, Complex unit gain bicyclic graphs with rank 2, 3 or 4. Linear Algebra Appl. 523(2017) 169--186.



\bibitem{LWZ} Yong Lu, Ligong Wang, Qiannan Zhou, Bicyclic oriented graphs with skew-rank 6, Appl. Math. Comput.  270(2015) 899--908.

\bibitem{LWZ1} Yong Lu, Ligong Wang, Qiannan Zhou, The rank of a signed graph in terms of the rank of its underlying graph, Linear Algebra Appl. 538(2018) 166--186.
\bibitem{LWZ2} Yong Lu, Ligong Wang, Qiannan Zhou, Skew-rank of an oriented graph in terms of the rank and dimension of cycle space of its underlying graph, Filomat 32(2018) 1303--1312.

\bibitem{LWZ3} Yong Lu, Ligong Wang, Qiannan Zhou, The rank of a complex unit gain graph in terms of the rank of its underlying graph. J. Comb.Optim. 38(2019)  570--588.

\bibitem{LUWU} Yong Lu, Jingwen Wu, Bounds for the rank of a complex unit gain graph in terms of its maximum degree, Linear Algebra Appl. 610(2021) 73--85.
 \bibitem{LUWU1}Yong Lu, Jingwen Wu, No signed graph with the nullity $\eta(G,\sigma)=|V(G)|-2m(G)+2c(G)-1$. Linear Algebra Appl. 615(2021), 175--193.

%\bibitem{MFANG} Xiaobin Ma, Xianwen Fang, An improved lower bound for the nullity of a graph in terms of matching number, Linear Multilinear Algebra. 2019. https://doi.org/10.1080/03081087.2019.1567672.

%\bibitem{MWTDAM} Xiaobin Ma,  Dein Wong, Fenglei Tian, Nullity of a graph in terms of the dimension of cycle space and the number of pendant vertices, Discrete Appl. Math. 215(2016) 171--176.


\bibitem{MWTLAA} Xiaobin Ma,  Dein Wong, Fenglei Tian, Skew-rank of an oriented graph in terms of matching number, Linear Algebra Appl. 495(2016) 242--255.



\bibitem{QHYG} Hui Qu, Guihai Yu, Bicyclic oriented graphs with skew-rank 2 or 4, Appl. Math. Comput.  258(2015) 182--191.

%\bibitem{QHYGFL}Hui Qu, Guihai Yu, Lihua Feng, More on the minimum skew-rank of graphs, Oper. Matrices. 9(2015) 311--324.


\bibitem{NR} N. Reff, Spectral properties of complex unit gain graphs. Linear Algebra Appl. 436(2012) 3165--3176.


\bibitem{RCZ} S. Rula, An Chang, Yirong Zheng, The extremal graphs with respect to their nullity, J. Inequal. Appl. 2016(2016) 71, 13 pp.



%\bibitem{SST} Yazhi Song, Xiaoqiu Song, Bit-Shun Tam, A Characterization of graphs $G$ with nullity $V(G)-2m(G)+2c(G)$, Linear Algebra Appl. 465(2015) 363--375.

%\bibitem{SSZ} Yazhi Song, Xiaoqiu Song, Mingcui Zhang, An upper bound for the nullity of a bipartite graph in terms of its maximum degree, Linear Multilinear Algebra. 64(2016) 1107--1112.

%\bibitem{TW} Fenglei Tian, Dein Wong, Relation between the skew energy of an oriented graph and its matching number. Discrete Appl. Math. 222(2017) 179--184.

\bibitem{LWANG} Long Wang, Characterization of graphs with given order, given size and given matching number that minimize nullity, Discrete Math. 339(2016) 1574--1582.

\bibitem{WW} Long Wang, Dein Wong, Bounds for the matching number, the edge charomatic numbber and the independence number of a graph in terms of rank, Discrete Appl. Math. 166(2014) 276--281.

\bibitem{WLMA} Long Wang, Nullity of a graph in terms of path cover number, Linear Multilinear Algebra. 69(2021) 1902--1908.

\bibitem{WGF} Yi Wang, Shi-Cai Gong, Yi-Zheng Fan, On the determinant of the Laplacian matrix of a complex unit gain graph, Discrete Math. 341(2018) 81--86.

\bibitem{WMT} Dein Wong, Xiaobin Ma, Fenglei Tian, Relation between the skew-rank of an oriented graph and the rank of its underlying graph, European J. Combin. 54(2016) 76--86.

\bibitem{WLM} Wei Wei, Shuchao Li, Hongping Ma, Bounds on the nullity, the $H$-rank and the Hermitian energy of a mixed graph, Linear Multilinear Algebra. 2019. https://doi.org/10.1080/03081087.2019.1679702.

\bibitem{XZWT}Feng Xu, Qi Zhou, Dein Wong, Fenglei Tian, Complex unit gain graphs of rank 2, Linear Algebra Appl. 597(2020) 155--169.

\bibitem{YQT} Guihai Yu, Hui Qu, Jianhua Tu, Inertia of complex unit gain graphs, Appl. Math. Comput.  265(2015) 619--629.

\bibitem{YFQ} Guihai Yu, Lihua Feng, Hui Qu, Signed graphs with small positive index of inertia, Electron. J. Linear Algebra 31(2016) 232--243.



\bibitem{ZaHe} S. Zaman, Xiaocong He, Relation between the inertia indices of a complex unit gain graph and those of its underlying graph, 2020. https://doi.org/10.1080/03081087.2020.1749224.

\bibitem{ZWS} Qi Zhou, Dein Wong, Dongqin Sun, An upper bound of the nullity of a graph in terms of order and maximum degree, Linear Algebra Appl. 555(2018) 314--320.



\end{thebibliography}

\documentclass[11pt]{article}
\usepackage{enumerate}
\usepackage{amssymb,a4wide,latexsym,makeidx,epsfig,fleqn}
\usepackage{amsthm}
\usepackage{amsmath}
\usepackage{enumerate}
\newtheorem{theorem}{Theorem}[section]
\newtheorem{remark}[theorem]{Remark}

\newtheorem{definition}[theorem]{Definition}
\newtheorem{lemma}[theorem]{Lemma}

\begin{document}
\textwidth 150mm \textheight 225mm
\title{Inertia indices of a complex unit gain graph in terms of matching number  \footnote{This work is supported by the National Natural Science Foundations of China (No. 11901253), the Natural Science Foundation for Colleges and Universities in Jiangsu Province of China (No. 19KJB110009), and the Science Foundation of Jiangsu Normal University (No.18XLRX021).}}
\author{{Yong Lu and Qi Wu\footnote{Corresponding author.}}\\
{\small  School of Mathematics and Statistics, Jiangsu Normal University,}\\ {\small  Xuzhou, Jiangsu 221116,
People's Republic
of China.}\\
{\small E-mails:luyong@jsnu.edu.cn, wuqimath@163.com}}

\date{}
\maketitle
\begin{center}
\begin{minipage}{120mm}
\vskip 0.3cm
\begin{center}
{\small {\bf Abstract}}
\end{center}
{\small A complex unit gain graph  is a triple
$\varphi=(G, \mathbb{T}, \varphi)$ (or $G^{\varphi}$ for short) consisting of a simple
graph $G$, as the underlying graph of $G^{\varphi}$, the set of unit
complex numbers $\mathbb{T}={z\in \mathbb{C}: |z| = 1}$ and a gain function
$\varphi: \overrightarrow{E}\rightarrow \mathbb{T}$ such that $\varphi(e_{i,j})=\varphi(e_{j,i}) ^{-1}$. Let $A(G^{\varphi})$ be adjacency matrix of $G^{\varphi}$. In this paper, we prove that
$$m(G)-c(G)\leq p(G^{\varphi})\leq m(G)+c(G),$$
$$m(G)-c(G)\leq n(G^{\varphi})\leq m(G)+c(G),$$
where $p(G^{\varphi})$, $n(G^{\varphi})$, $m(G)$ and $c(G)$ are the number of positive eigenvalues of $A(G^{\varphi})$, the number of negative eigenvalues of $A(G^{\varphi})$, the matching number and the cyclomatic number of $G$, respectively. Furthermore, we characterize
the graphs which attain the upper bounds and the lower
bounds, respectively.

\vskip 0.1in \noindent {\bf Key Words}: \ Complex unit gain graph; Inertia index; Matching number; Cyclomatic number. \vskip
0.1in \noindent {\bf AMS Subject Classification (2010)}: \ 05C35; 05C50. }
\end{minipage}
\end{center}

\section{Introduction }
 All graphs in this paper are simple graphs, without multiedges and loops. Let $G=(V(G), E(G))$
 be a simple graph and $V(G)=\{v_{1},v_{2},\ldots,v_{n}\}$, $E(G)$ be the vertex set and edge set of $G$, respectively. Denote by $e_{ij}$  the oriented edge from  $v_{i}$ to $v_{j}$. Let $\overrightarrow{E}$ be the oriented edge set obtained from $E(G)$, and $\{e_{ij},e_{ji}\}\in \overrightarrow{E}$. Denote by $N_{G}(x)$  the \emph{neighbor set} of a vertex $x\in V(G)$, and $d_{G}(x)=|N_{G}(x)|$  the \emph{degree} of $x$. In particular, $x$ is called a \emph{pendant} vertex of $G$ if $d_{G}(x)=1$. The neighbour vertex of a pendant vertex is called a \emph{quasi-pendant} vertex in graph $G$. We use $P_{n}$, $C_{n}$ to denote a \emph{path}, a \emph{cycle} of order $n$, respectively. An induced subgraph $C_{p}$ of a graph $G$ is called a \emph{pendant cycle} if $C_{p}$ is a cycle and has a unique vertex of degree 3 in $G$. Denote by $F(G)$ the edge set that every edge in $F(G)$  has an endpoint on some cycle of $G$ and other endpoint outside the cycle of $G$.

  Let $c(G)=|E(G)|-|V(G)|+\omega(G)$ be the \emph{cyclomatic number} of a graph $G$, where $\omega(G)$ is the number of connected components of $G$. Two distinct edges in a graph $G$ are \emph{independent} if they do not have common end-vertex in $G$. A set of pairwise independent edges of $G$ is called a \emph{matching} of $G$. A matching with the maximum cardinality is a \emph{maximum matching} of $G$, denote by $M(G)$. The \emph{matching number} of $G$, denoted by $m(G)$, is the cardinality of a maximum matching of $G$. A maximum matching $M(G)$ of $G$ \emph{saturates} $v$ of $G$ if $v$ is an endpoint of an edge in $M(G)$. Two graphs are called \emph{vertex-disjoint} if they have no common vertices.
  Let $G$ be a graph with pairwise vertex-disjoint cycles. $T_{G}$ is an acyclic graph obtained from $G$ by contracting each cycle of $G$ into a vertex, called a \emph{cyclic vertex}. Denote by $O(G)$  the set of vertices in cycles of $G$.

 Denote by $\mathfrak{G}$ a class of graphs that every graph $G\in\mathfrak{G}$ hold the following properties: (1) $G$ contains at least one cycle but  is not the disjoint union the disjoint cycles and/or trees, (2) any two cycles of $G$ share no common vertices if $G$ contains more than one cycle.

The \emph{adjacency matrix}  of $G$, denote by $A(G)$ is the symmetric $n\times n$ matrix with entries  $a_{ij}=1$ if and only if $v_{i}v_{j}\in E(G)$ and 0 elsewhere. Denote by $r(G)$ the \emph{rank} of $G$, which is the rank of $A(G)$. The multiplicity of 0 as an eigenvalue of $A(G)$, denoted by $\eta(G)$, is called the \emph{nullity} of $G$. The \emph{positive inertia index}(resp.the \emph{negative inertia index}), denoted by $p(G)$(resp.$n(G)$) is the number of positive eigenvalues(resp.negative eigenvalues) of $A(G)$. For a  simple graph of order $n$, $r(G)=p(G)+n(G)=n-\eta(G)$.

Denote by $\Phi=(G,\mathbb{T},\varphi)$ a complex unit gain graph, where $G$ is the \emph{underlying graph} of $\Phi$, $\mathbb{T}=\{z\in C: |z|=1\}$ is the \emph{circle group}, and   $\varphi:\overrightarrow{E}\rightarrow \mathbb{T}$, where $\varphi(e_{ij})=\varphi(e_{ji})^{-1}=\overline{\varphi(e_{ji})}$. For convenience, $\varphi(e_{ij})$ is also written as $\varphi_{v_{i}v_{j}}$ for $v_{i}v_{j}\in E(\Phi)$.  Denote by $G^{\varphi}$  a complex unit gain  graph $\Phi=(G,\mathbb{T},\varphi)$ for convenience. The \emph{adjacency matrix} of  $G^{\varphi}$ is the  Hermitian matrix $A(G^{\varphi})=(a_{ij})_{n\times n}$, where $a_{ij}=\varphi(e_{ij})$ if $v_{i}v_{j}\in E(G)$, and $a_{ij}=0$ otherwise. The rank $r(G^{\varphi})$ of $A(G^{\varphi})$ is called the \emph{rank}  of $G^{\varphi}$.The \emph{positive inertia index}(resp.the \emph{negative inertia index}), denoted by $p(G^{\varphi})$(resp.$n(G^{\varphi})$) is the number of positive eigenvalues(resp.negative eigenvalues) of $A(G^{\varphi})$. For a complex unit gain graph $G^{\varphi}$, the matching number, the cyclomatic number, pendant vertex and quasi-pendant vertex of $G^{\varphi}$ are defined to be the matching number, the cyclomatic number, pendant vertex and quasi-pendant vertex of its underlying graph, respectively.

 We write $G^{\varphi}-v$ ($v\in V(G^{\varphi})$) for the \emph{induced subgraph} obtained from $G^{\varphi}$ by deleting $v$ and all its incident edges. Let $V(G_{1}^{\varphi})\subset V(G^{\varphi})$, denote by $G^{\varphi}-G^{\varphi}_{1}$ the induced subgraph obtained from $G^{\varphi}$ by deleting all vertices of $G^{\varphi}_{1}$ and all incident edges. For $V(G^{\varphi}_{2})\subset V(G^{\varphi})$ and $v\notin G^{\varphi}_{2}$, denote by $G^{\varphi}_{2}+v$, the induced subgraph of $G^{\varphi}$ with vertex set $V(G^{\varphi}_{2})\cup \{v\}$. Let $G^{\varphi}$, $H^{\varphi}$ be two complex unit gain graphs, denote by $G^{\varphi}\cup H^{\varphi}$ the disjoin union of $G^{\varphi}$ and $H^{\varphi}$.

 Collatz et al.\cite{CS} had wanted to obtain all graphs of order $n$ with $r(G)<n$. Until today, this problem is also unsolved.
In mathematics, the rank (or nullity, inertia index) of a graph has attracted a lot of researchers' attention, they focus on the  bounds for  the rank (or nullity, inertia index) of a simple graph $G$ \cite{FYZW, FHLL, LGUO, RCZ, LWANG, WW, WLMA,ZWS}, a signed graph \cite{FDD,FWW, HHL,  LWZ1,LUWU1, YFQ}, an oriented graph \cite{LXYG, LWZ, LWZ2, MWTLAA, QHYG, WMT} and a mixed graph \cite{CHL,WLM} and so on.

For a complex unit gain graph $G^{\varphi}$, Reff \cite{NR}  defined the adjacency, incidence and Laplacian matrices of a complex unit gain graph. Some eigenvalue bounds for the adjacency and Laplacian matrices were present. Yu et al.\cite{YQT} give the inertia of some complex unit gain graph. Lu et al.\cite{LWX} characterized all the complex unit gain bicyclic graphs $G^{\varphi}$ with $r(G^{\varphi})=2,3,4$. Lu et al.\cite{LWZ3} obtained relation between the rank of a complex unit gain graph in terms of the rank of its underlying graph. Wang et al.\cite{WGF} obtained the  determinant of the Laplacian matrix of a complex unit gain graph. Xu et al.\cite{XZWT} characterized all the complex unit gain graphs of rank 2.  He et al.\cite{HHY} obtained the  bounds for the rank of a complex unit gain graph in terms of the independence number. Zaman and  He \cite{ZaHe} obtained the relation between the inertia indices of a complex unit gain graph and those of its underlying graph.  Lu and Wu \cite{LUWU} obtained the bounds for the rank of a complex unit gain graph in terms of its maximum degree.

In 2020, Li, Wang\cite{LiW} and He et al.\cite{HHD} obtained the rank of a complex unit graph in terms of the matching number, respectively. They proved that $$2m(G)-2c(G)\leq r(G^{\varphi})\leq 2m(G)+c(G).$$  All corresponding extremal graphs are characterized by them. Motivated by their results, in this paper, we will prove that $$m(G)-c(G)\leq p(G^{\varphi})\leq m(G)+c(G),$$
$$m(G)-c(G)\leq n(G^{\varphi})\leq m(G)+c(G).$$  All corresponding extremal graphs are characterized.

 In Section 2, we give and prove some lemmas about complex unit gain graphs. In Section 3, we characterize the relations between the inertia indices of a complex unit gain graph and its matching number.

\section{Preliminaries}
\quad For a simple graph $G$, there has some lemmas.
\noindent\begin{lemma}\label{le:2.1}\cite{CHL}
Let $G$ be a simple graph. Then $m(G)-1\leq m(G-v)\leq m(G)$ for any vertex $v\in V(G)$.
\end{lemma}
\noindent\begin{lemma}\label{le:2.2}\cite{HHD}
Let $G$ be a graph obtained by joining a vertex of an even cycle $C$ by an edge to a vertex of a connected graph $H$. Then $m(G)=m(C)+m(H)$.
\end{lemma}
\noindent\begin{lemma}\label{le:2.3}\cite{HHD}
Let $x$ be a pendant vertex of a graph G and $y$ be the neighbour of $x$. Then $m(G)=m(G-y)+1=m(G-x-y)+1$.
\end{lemma}
\noindent\begin{lemma}\label{le:2.4}\cite{FYZW}
Let $G$ be a graph with at least one cycle. Suppose that all cycles of $G$
are pairwise vertex-disjoint and each cycle is odd, then $m(T_{G})=m(G-O(G))$ if and only if there exists a maximum matching $M(G)$ of $G$ such that $M(G)\cap F(G)=\varnothing$.
\end{lemma}
\noindent\begin{lemma}\label{le:2.13}\cite{FYZW}
Let $G\in \mathfrak{G}$. If $m(T_{G})=m(G-O(G))$, then
$G$ contains at least one pendant vertex, and any quasi-pendant vertex of
$G$ lies outside of cycles.
\end{lemma}
\noindent\begin{lemma}\label{le:2.6}\cite{WMT}
Let $G$ be a graph with $x\in V (G)$. Then
\begin{enumerate}[(a)]

\item $c(G)=c(G-x)$ if $x$ lies outside any cycle of $G$;

\item $c(G-x)\leq c(G)-1$ if $x$ lies on a cycle of $G$;

\item $c(G-x)\leq c(G)-2$ if $x$ is a common vertex of distinct cycles of $G$.
\end{enumerate}
\end{lemma}
For a complex unit gain graph, we have the following definition and lemmas.

\noindent\begin{definition}\label{de:2.7}\cite{LWX}
Let $C_{n}^{\varphi} (n\geq3)$ be a complex unit gain cycle, denote by $$\varphi(C_{n})=\varphi_{v_{1}v_{2}}\varphi_{v_{2}v_{3}}\cdots \varphi_{v_{n-1}v_{n}}\varphi_{v_{n}v_{1}}.$$ Then $C_{n}^{\varphi}$ is said to be:

\begin{displaymath}
\left\{\
        \begin{array}{ll}
          \rm Type~A,&  \emph{if}~\varphi(C_{n})=(-1)^{n/2}~\emph{and}~n~\emph{is~even},\\
          \rm Type~B,& \emph{if}~\varphi(C_{n})\neq(-1)^{n/2}~\emph{and}~n~\emph{is~even},\\
          \rm Type~C,& \emph{if}~Re\left((-1)^{{(n-1)}/{2}}\varphi(C_{n})\right)>0~\emph{and}~n~\emph{is~odd},\\
          \rm Type~D,& \emph{if}~Re\left((-1)^{{(n-1)}/{2}}\varphi(C_{n})\right)<0~\emph{and}~n~\emph{is~odd},\\
          \rm Type~E,& \emph{if}~Re\left((-1)^{{(n-1)}/{2}}\varphi(C_{n})\right)=0~\emph{and}~n~\emph{is~odd}.
        \end{array}
      \right.
\end{displaymath}
\end{definition}
where $Re(\cdot)$ is the real part of a complex number.
\noindent\begin{lemma}\label{le:2.8}\cite{YQT}
Let $C_{n}^{\varphi}$ be a complex unit gain cycle of order $n$. Then

\begin{displaymath}
(p(C_{n}^{\varphi}),n(C_{n}^{\varphi}))=\left\{\
        \begin{array}{ll}
          (\frac{n-2}{2},\frac{n-2}{2}),& \emph{if}~C_{n}^{\varphi}\rm~is~of~Type~A,\\
          (\frac{n}{2},\frac{n}{2}),& \emph{if}~C_{n}^{\varphi}\rm~is~of~Type~B,\\
          (\frac{n+1}{2},\frac{n-1}{2}),& \emph{if}~C_{n}^{\varphi}\rm~is~of~Type~C,\\
          (\frac{n-1}{2},\frac{n+1}{2}),& \emph{if}~C_{n}^{\varphi}\rm~is~of~Type~D,\\
          (\frac{n-1}{2},\frac{n-1}{2}),& \emph{if}~C_{n}^{\varphi}\rm~is~of~Type~E.
        \end{array}
      \right.
\end{displaymath}
\end{lemma}

\noindent\begin{lemma}\label{le:2.9}\cite{YQT}
 Let $G^{\varphi}$ be a complex unit gain graph.
\begin{enumerate}[(a)]
\item Let $H^{\varphi}$ be an induced subgraph of $G^{\varphi}$. Then $p(H^{\varphi})\leq p(G^{\varphi})$ and $n(H^{\varphi})\leq n(G^{\varphi})$.
\item Let $G_{1}^{\varphi},G_{2}^{\varphi},\cdots,G_{t}^{\varphi}$ be the connected components of $G^{\varphi}$. Then $p(G^{\varphi})=\sum_{i=1}^{t} p(G_{i}^{\varphi})$ and $n(G^{\varphi})=\sum_{i=1}^{t} n(G_{i}^{\varphi})$.
\item $p(G^{\varphi})=0$ ($n(G^{\varphi})=0$) if and only if $G^{\varphi}$ is a graph without edges.
\end{enumerate}
\end{lemma}

\noindent\begin{lemma}\label{le:2.10}\cite{YQT}
Let $T^{\varphi}$ be an acyclic complex unit gain graph. Then $p(T^{\varphi})=n(T^{\varphi})=m(T)$.
\end{lemma}
\noindent\begin{lemma}\label{le:2.11}\cite{YQT}
Let $G^{\varphi}$ be a complex unit gain graph that has a pendant vertex $u$ and $v$ is the unique neighbour of $u$. Then $p(G^{\varphi})=p(G^{\varphi}-v)+1=p(G^{\varphi}-u-v)+1,n(G^{\varphi})=n(G^{\varphi}-v)+1=n(G^{\varphi}-u-v)+1$.
\end{lemma}
\noindent\begin{lemma}\label{le:2.12}\cite{ZaHe}
Let $G^{\varphi}$ be a complex unit gain graph and $u$ is a vertex of $G^{\varphi}$. Then $p(G^{\varphi})-1\leq p(G^{\varphi}-u)\leq p(G^{\varphi}), n(G^{\varphi})-1\leq n(G^{\varphi}-u)\leq n(G^{\varphi})$.
\end{lemma}

\section{Bounds for the inertia indices of $G^{\varphi}$}
\quad In this section, we will  obtain some bounds for the inertia indices of a complex unit gain graph in terms of its matching number.

\noindent\begin{theorem}\label{th:3.1}
Let $G^{\varphi}$ be a  connected complex unit gain graph. Then
\begin{enumerate}[(a)]
\item $m(G)-c(G)\leq p(G^{\varphi})\leq m(G)+c(G)$.
\item $m(G)-c(G)\leq n(G^{\varphi})\leq m(G)+c(G)$.
\end{enumerate}
\end{theorem}
\noindent\textbf{Proof.}
we get the inequalities above by induction on $c(G)$. If $c(G)=0$, then $G^{\varphi}$ is a tree, $p(G^{\varphi})=n(G^{\varphi})=m(G)$ holds by Lemma \ref{le:2.10}.
Now we assume that assertion holds for every connected complex unit gain graph that the cyclomatic number is less than $c(G)$. Let $v$ be a vertex on some cycle in $G^{\varphi}$. Let $H_{1},\cdots,H_{k}$ be the connected components of $G-v$. By Lemma \ref{le:2.6}(b), we have
$$\sum_{i=1}^{k} c(H_{i})=c(G-v)\leq c(G)-1.$$
Combining with  Lemmas \ref{le:2.1}, \ref{le:2.9}(b) and \ref{le:2.12}, we have
$$\sum_{i=1}^{k} p(H_{i}^{\varphi})=p(G^{\varphi}-v)\leq p(G^{\varphi})\leq p(G^{\varphi}-v)+1$$
and
$$m(G)-1\leq m(G-v)=\sum_{i=1}^{k} m(H_{i})\leq m(G).$$

Because $c(G-v)\leq c(G)-1$, by the induction hypothesis, for each $i\in\{1,2,\ldots,k\}$,
$$m(H_{i})-c(H_{i})\leq p(H_{i}^{\varphi})\leq m(H_{i})+c(H_{i})$$

Hence,
\begin{align*}
&p(G^{\varphi})\geq p(G^{\varphi}-v)\\
&~~~~~~~~=\sum_{i=1}^{k} p(H_{i}^{\varphi})\\
&~~~~~~~~\geq \sum_{i=1}^{k} [m(H_{i})-c(H_{i})]\\
&~~~~~~~~\geq m(G)-1-(c(G)-1)\\
&~~~~~~~~=m(G)-c(G).
\end{align*}
and
\begin{align*}
&p(G^{\varphi})\leq p(G^{\varphi}-v)+1\\
&~~~~~~~~=\sum_{i=1}^{k} p(H_{i}^{\varphi})+1\\
&~~~~~~~~~\leq \sum_{i=1}^{k} [m(H_{i})+c(H_{i})]+1\\
&~~~~~~~~~\leq m(G)+c(G)-1+1\\
&~~~~~~~~~=m(G)+c(G).
\end{align*}
That is $$m(G)-c(G)\leq p(G^{\varphi})\leq m(G)+c(G).$$
Similarly, we can obtain that
$$m(G)-c(G)\leq n(G^{\varphi})\leq m(G)+c(G).$$

This completes the proof.  \quad $\square$

For convenience, we call $G^{\varphi}$ to be $p$-lower optimal(resp.,$p$-upper optimal) if $p(G^{\varphi})$ obtains the lower bound (upper bound) in Theorem \ref{th:3.1}. Similarly, $G^{\varphi}$ is $n$-lower optimal(resp.,$n$-upper optimal) if $n(G^{\varphi})$ obtains the lower bound (upper bound) in Theorem \ref{th:3.1}.

\noindent\begin{lemma}\label{le:3.3}
Let $G^{\varphi}$ be a complex unit gain graph and $H_{1}^{\varphi},H_{2}^{\varphi},\cdots,H_{k}^{\varphi}$ are its connected components. Then
\begin{enumerate}[(a)]
  \item $G^{\varphi}$ is $p$-lower optimal if and only if for each $i\in\{1,2,\ldots,k\}$, $H_{i}^{\varphi}$ is $p$-lower optimal.
  \item $G^{\varphi}$ is $p$-upper optimal if and only if for each $i\in\{1,2,\ldots,k\}$, $H_{i}^{\varphi}$ is $p$-upper optimal.
\end{enumerate}

\end{lemma}
\noindent\textbf{Proof.} At first, we will prove the (a) of this lemma.

\textbf{Necessity:} Assume that the conclusion is not true, without loss of generality, we suppose that $H_{1}^{\varphi}$ is not $p$-lower optimal. By Theorem \ref{th:3.1},
$$p(H_{1}^{\varphi})>m(H_{1})-c(H_{1}),$$
and for each $i\in\{2,3,\ldots,k\}$, we have
$$p(H_{i}^{\varphi})\geq m(H_{i})-c(H_{i}).$$

By Lemma \ref{le:2.9}(b), we have $$p(G^{\varphi})=\sum_{i=1}^{k} p(H_{i}^{\varphi})>\sum_{i=1}^{k} [m(H_{i})-c(H_{i})]=m(G)-c(G),$$
a contradiction.

\textbf{Sufficiency:} For each $i\in\{1,2,\ldots,k\}$, $H_{i}^{\varphi}$ is $p$-lower optimal, so we have
$$p(H_{i}^{\varphi})=m(H_{i})-c(H_{i}).$$

By Lemma \ref{le:2.9}(b), we have $$p(G^{\varphi})=\sum_{i=1}^{k} p(H_{i}^{\varphi})=\sum_{i=1}^{k} [m(H_{i})-c(H_{i})]=m(G)-c(G).$$

Using the same method, we can prove the (b) of this lemma.

This complete the proof.   \quad $\square$

\noindent\begin{lemma}\label{le:3.4}
Let $G^{\varphi}$ be a complex unit gain graph and a vertex $v$ lies on some cycle of $G^{\varphi}$.
Then the following results are established:
\begin{enumerate}[(a)]
\item If $G^{\varphi}$ is $p$-lower optimal, then $p(G^{\varphi})=p(G^{\varphi}-v)$, $p(G^{\varphi}-v)=m(G-v)-c(G-v)$, $c(G)=c(G-v)+1$, $m(G)=m(G-v)+1$;
\item If $G^{\varphi}$ is $p$-upper optimal, then $p(G^{\varphi})=p(G^{\varphi}-v)+1$, $p(G^{\varphi}-v)=m(G-v)+c(G-v)$, $c(G)=c(G-v)+1$, $m(G)= m(G-v)$;
\item If $G^{\varphi}$ is $p$-lower optimal (or upper optimal), then $v$ lies on just one cycle of $G^{\varphi}$ and $v$ is not a quasi-pendant vertex.
\end{enumerate}
\end{lemma}
\noindent\textbf{Proof.}
By the proof of Theorem \ref{th:3.1}, we can obtain the (a) and (b). For (c), when $G^{\varphi}$ is $p$-lower optimal. If $v$ lie on at least two cycles of $G^{\varphi}$, by Lemma \ref{le:2.6}(c),
$$c(G-v)\leq c(G)-2,$$ which contradicts (a). If $v$ is a quasi-pendant vertex of $G^{\varphi}$, then by Lemma \ref{le:2.11},
$$p(G^{\varphi}-v)=p(G^{\varphi})-1,$$ which contradicts (a).

When $G^{\varphi}$ is $p$-upper optimal. If $v$ lie on at least two cycles of $G^{\varphi}$, by Lemma \ref{le:2.6}(c),
$$c(G-v)\leq c(G)-2,$$ which contradicts (b). If $v$ is a quasi-pendant vertex of $G^{\varphi}$, then by Lemma \ref{le:2.3},
$$m(G)=m(G-v)+1,$$ which contradicts (b).
 So the assertion (c) is  hold.   \quad $\square$

\noindent\begin{lemma}\label{le:3.5}
Let  $G^{\varphi}$ be a complex unit gain graph which contains a pendent vertex $x$ with its unique neighbour $y$. Let $H^{\varphi}=G^{\varphi}-x-y$. If $G^{\varphi}$ is $p$-lower optimal (upper optimal), then $H^{\varphi}$ is also $p$-lower optimal (upper optimal).
\end{lemma}
\noindent\textbf{Proof.}
If $G^{\varphi}$ is $p$-lower optimal, we have
$$p(G^{\varphi})=m(G)-c(G).$$

By Lemmas \ref{le:2.3}, \ref{le:2.6}(a) and \ref{le:2.11}, we have $$p(H^{\varphi})=p(G^{\varphi})-1=m(G)-c(G)-1=m(H)+1-c(H)-1=m(H)-c(H),$$
so $H^{\varphi}$ is also $p$-lower optimal.

Using the same method, we can get that if $G^{\varphi}$ is $p$-upper optimal, then $H^{\varphi}$ is $p$-upper optimal.
  \quad $\square$

\noindent\begin{lemma}\label{le:3.6}
Let $G^{\varphi}$ be a complex unit gain unicyclic graph and $C_{q}^{\varphi}$ is unique  cycle of $G^{\varphi}$.
\begin{enumerate}[(a)]
  \item If $G^{\varphi}$ is $p$-lower optimal, then $C_{q}^{\varphi}$ is of Type A.
  \item If $G^{\varphi}$ is $p$-upper optimal, then $C_{q}^{\varphi}$ is of Type C.
\end{enumerate}
\end{lemma}
\noindent\textbf{Proof.} At first, we will prove the (a) of this lemma. We shall apply induction on the order of $T_{G}$.
 If $|V(T_{G})|=1$, then $G^{\varphi}$ is $C_{q}^{\varphi}$.  By Lemma \ref{le:2.8}, we can get that $C_{q}^{\varphi}$ is of Type A.
 Now suppose that $|V(T_{G})|\geq2$, if $G^{\varphi}$ has no pendent vertices, $G^{\varphi}$ is the union of $C_{q}^{\varphi}$ and some isolated vertices, by Lemmas \ref{le:2.8} and \ref{le:2.9}(b), $C_{q}^{\varphi}$ is of Type A. If $G^{\varphi}$ has a pendent vertex, say $u$ and $v$ is its neighbour vertex in $G^{\varphi}$. By Lemma \ref{le:3.4}(c), $v$ is not on $C_{q}^{\varphi}$, so $|V(T_{G})|\neq 2$, $|V(T_{G})|\geq 3$. 
Then by Lemma \ref{le:3.5}, $G^{\varphi}-u-v$ is $p$-lower optimal and $C_{q}^{\varphi}$ is its unique cycle, by induction hypothesis, $C_{q}^{\varphi}$ is of Type A.

Using the same method, we can obtain the (b) of this lemma.
\quad $\square$

\noindent\begin{lemma}\label{le:3.7}
Let $H^{\varphi}$ be a complex unit gain graph with any two cycles (if any) share no common vertices. Denote by $G^{\varphi}$ the graph obtained from adding an edge between a vertex $u$ of a  cycle $C_{q}^{\varphi}$ and a vertex $v$ of $H^{\varphi}$. If $G^{\varphi}$ is $p$-lower optimal, then
\begin{enumerate}[(a)]
\item Every  cycle of $G^{\varphi}$ is of Type $A$;
\item The edge $uv$ does not belong to any maximum matching of $G$;
\item Each maximum matching of $H$ saturates $v$;
\item $m(H+u)=m(H)$; 
\item $H^{\varphi}$ is $p$-lower optimal;
\item Let $K^{\varphi}$ be the induced  subgraph of $G^{\varphi}$ with vertex set
$V(H)\cup{x}$. Then $K^{\varphi}$ is also $p$-lower optimal;
\end{enumerate}
\end{lemma}
\noindent\textbf{Proof.}
We use induction on $c(G)$, since $G^{\varphi}$ has a cycle $C_{q}^{\varphi}$, $c(G)\geq 1$.
If $H^{\varphi}$ contains no cycle, by Lemma \ref{le:3.6}(a), we have that $C_{q}^{\varphi}$ is of Type A.

If $H^{\varphi}$ contains at least one cycle. Let $x$ be a vertex lying on some cycle of $H^{\varphi}$ and $G_{0}^{\varphi}=G^{\varphi}-x$. By Lemma \ref{le:3.4}(a), we have $G_{0}^{\varphi}$ is $p$-lower optimal.

By induction hypothesis, one has that each cycle in $G_{0}^{\varphi}$, including $C_{q}^{\varphi}$ is of Type A. By a
similar discussion as for $G^{\varphi}-u$, we can show that each cycle in $H^{\varphi}$ is of Type A.
This completes the proof of (a).

Suppose on the contrary that $uv$ belongs to a maximum matching
$M$ of $G$. By (a),  $C_{q}^{\varphi}$ is an even cycle, so there exists a vertex $w$ lying on $C_{q}^{\varphi}$
such that $w$ is not saturated by $M$. Thus we have
$$m(G)=m(G-w),$$ a contradiction to Lemma \ref{le:3.4}(a). This completes the proof of (b).

By Lemma \ref{le:2.2}, we have $$m(G)=m(H)+m(C_{q}).$$ Let $M_{1}$ be the maximum
matching of $C_{q}$. Suppose on the contrary that there exists a maximum matching $M_{2}$ of $H$ fails to saturate $v$, so $M_{1} \cup M_{2}$ is a maximum matching of $G$. Then we obtain a maximum matching $M^{'} \cup M_{2}$ of $G$ which contains $uv$, where $M^{'}$ is
obtained from $M_{1}$ by replacing the edge in $M_{1}$ which saturates $u$ with $uv$, a contradiction
to (b). This completes the proof of (c).

By (c), we can obtain the (d) of this lemma.

 By Lemma \ref{le:3.4}(a), $G^{\varphi}-u$ is $p$-lower optimal. Then (e) immediately follows
from Lemma \ref{le:3.3}(a).

Suppose that $C_{q}=uu_{2}u_{3} \cdots u_{2s}u$. Since $G^{\varphi}$ is $p$-lower optimal, by Lemma \ref{le:3.4}(a), one has that $G_{1}^{\varphi}=G^{\varphi}-u_{2}$ is also $p$-lower optimal. Obviously, $u_{3}$ and $u_{4}$ are
pendant vertex and quasi-pendant vertex of $G_{1}^{\varphi}$, respectively. By Lemma \ref{le:3.5}, one has
that $G_{2}^{\varphi}=G_{1}^{\varphi}-u_{3}-u_{4}$ is also $p$-lower optimal. Repeating such process (deleting
a pendant vertex and a quasi-pendant vertex), after $s-1$ steps, the result graph is
$$G^{\varphi}-u_{2}-u_{3}-\cdots-u_{2s}=H^{\varphi}+u=K^{\varphi}.$$ By Lemma \ref{le:3.5}, $K^{\varphi}$ is also $p$-lower optimal.
   \quad $\square$

\noindent\begin{lemma}\label{le:3.8}
Let $G^{\varphi}$ be a connected complex unit gain graph. If $G^{\varphi}$ is $p$-lower optimal,
then there exists a maximum matching $M$ of $G$ such that $M \cap F(G)=\varnothing$. Moreover, $m(G)=m(G-O(G))+\sum_{C\subseteq G} m(C)$, where $C$ goes through all cycles of $G$.
\end{lemma}
\noindent\textbf{Proof.}
We shall apply induction on the order of $G^{\varphi}$. If $G^{\varphi}$ is an isolated vertex, the result holds trivially. Suppose $G^{\varphi}$ contains at least two vertices.
If $G^{\varphi}$ has a pendant vertex $u$, and $v$ is the unique neighbor of $u$. Let $H^{\varphi}=G^{\varphi}-u-v$
and $H_{1}^{\varphi},H_{2}^{\varphi},\cdots,H_{t}^{\varphi}$ be all connected components of $H^{\varphi}$. By Lemmas \ref{le:3.4}(c) and \ref{le:3.5}, $v$ is not on the cycle of $G^{\varphi}$ and $H^{\varphi}$ is $p$-lower optimal. Then by Lemma \ref{le:3.3}(a), we have $H_{i}^{\varphi}$ is also $p$-lower optimal for each $i\in\{1,2,\ldots,t\}$.

By induction hypothesis, for $H_{i}^{\varphi}$ there exists a maximum matching $M_{i}$ of $H_{i}$ such that $M_{i}\cap F(H_{i})=\varnothing$ for each $i\in \{1,2,\ldots,t\}$. Let $$M=(\cup_{i=1}^{t} M_{i} )\cup\{uv\}.$$ We can obtain  that $M$ is a
maximum matching of $G$ which satisfies $$M\cap F(G)=\varnothing.$$

If $G^{\varphi}$ has no pendant vertices, then $G^{\varphi}$ contains a pendant complex unit gain cycle, says $C^{\varphi}$. Let $$G_{1}^{\varphi}=G^{\varphi}-C^{\varphi}. $$ By Lemmas \ref{le:3.7}(a) and (e), it follows that $G_{1}^{\varphi}$ is $p$-lower optimal and the order of cycle $C^{\varphi}$ is even. Applying the induction on $G_{1}^{\varphi}$, there exists a maximum matching $M_{0}$ of $G_{1}$ such that $$M_{0} \cap F(G_{1})=\varnothing.$$

Let $M_{1}$ be a maximum matching of $C$. By Lemma \ref{le:2.2},  $M=M_{0}\cup M_{1}$
is a maximum matching of $G$ satisfying $$M\cap F(G)=\varnothing.$$  Moreover, we can get that
$$m(G)=m(G-O(G))+\sum_{C\subseteq G}m(C),$$ where $C$ goes through all cycles of $G$.

This completes the proof.   \quad $\square$
\noindent\begin{theorem}\label{th:3.9}
Let $G^{\varphi}$ be a connected complex unit gain graph. Then $G^{\varphi}$ is $p$-lower optimal if and only if the following
three conditions all hold:
\begin{enumerate}[(a)]
\item Any two cycles of $G^{\varphi}$  share no common vertices;
\item Each cycle of $G^{\varphi}$ is of Type A;
\item $m(T_{G})= m(G-O(G))$.
\end{enumerate}
\end{theorem}
\noindent\textbf{Proof.}
\textbf{Sufficiency:} We shall apply induction on the order of $G^{\varphi}$. If $G^{\varphi}$ is a complex unit gain
tree or single cycle of Type A,  the result follows from Lemmas \ref{le:2.8} and
\ref{le:2.10}. If not, since $m(T_{G})=m(G-O(G))$, we have that $G^{\varphi}$ contains at least one pendant vertex, say $x$. Let $y$ be the unique neighbour vertex of $x$ in $G^{\varphi}$. By Lemma \ref{le:2.13}, $y$ lies outside any cycle of $G^{\varphi}$.

Let $H^{\varphi}=G^{\varphi}-x-y$ and $H_{1}^{\varphi},H_{2}^{\varphi},\cdots,H_{t}^{\varphi}$ be all connected components
of $H^{\varphi}$.

Since $x,y$ are both outside the cycle of $G^{\varphi}$, then $$T_{H}=T_{G}-x-y$$ and
$$H-O(H)=G-O(G)-x-y.$$ By condition(c) and Lemma \ref{le:2.3}, we have
\begin{align*}
&m(T_{G})=\sum_{i=1}^{t} m(T_{H_{i}})+1\\
 &~~~~~~~~~=m(G-O(G))\\
 &~~~~~~~~~=\sum_{i=1}^{t} m(H_{i}-O(H_{i}))+1.
 \end{align*}

Therefore $$\sum_{i=1}^{t} m(T_{H_{i}})=\sum_{i=1}^{t} m(H_{i}-O(H_{i})).$$ We observe that for each $i\in\{1,2,\ldots,t\}$, $$m(T_{H_{i}})\geq m(H_{i}-O(H_{i})),$$ so we obtain $$m(T_{H_{i}})=m(H_{i}-O(H_{i}))$$ for each $i\in\{1,2,\ldots,t\}$. Therefore, $H_{i}^{\varphi}$ satisfies (a)-(c) for each $i\in \{1,2,\ldots,t\}$.

By induction hypothesis, for each $i\in \{1,2,\ldots,t\}$,
$$p(H_{i}^{\varphi})=m(H_{i})-c(H_{i}).$$
Then by Lemmas \ref{le:2.3}, \ref{le:2.9} and \ref{le:2.11},
\begin{align*}
&p(G^{\varphi})=p(H^{\varphi})+1\\
 &~~~~~~~~=\sum_{i=1}^{t} p(H_{i}^{\varphi})+1\\
 &~~~~~~~~=\sum_{i=1}^{t} [m(H_{i})-c(H_{i})]+1\\
 &~~~~~~~~=m(H)-c(H)+1\\
 &~~~~~~~~=m(G)-c(G).
 \end{align*}

\textbf{Necessity:} Let $G^{\varphi}$ be a complex unit gain graph such that $$p(G^{\varphi})=m(G)-c(G).$$ If $G^{\varphi}$ is a complex unit gain tree, $G^{\varphi}$ clearly satisfies (a)-(c) of this theorem. Assume that $G^{\varphi}$ has at least one complex unit gain cycle. The
assertion (a) follows from Lemma \ref{le:3.4}(c).

For (b), if $c(G)=1$, the result holds by Lemma \ref{le:3.6}. Now assume $c(G)=k$, where $k\geq 2$. If there exists a cycle, say $C_{1}^{\varphi}$, which is not of Type A, then by deleting an arbitrary vertex of each complex unit gain cycle of $G^{\varphi}$ except for
$C_{1}^{\varphi}$, we get a complex unit gain graph $H^{\varphi}$ with $c(H)=1$ and by Lemma \ref{le:3.6}, $$p(H^{\varphi})\geq m(H).$$ By Lemmas \ref{le:2.1} and \ref{le:2.9}(a),
\begin{align*}
&p(G^{\varphi})\geq p(H^{\varphi})\\
 &~~~~~~~~\geq m(H)\\
 &~~~~~~~~\geq m(G)-(k-1)\\
 &~~~~~~~~=m(G)-c(G)+1,
\end{align*}
a contradiction.

We prove the assertion (c) by the induction on the order of $G^{\varphi}$. If $G^{\varphi}$ is a complex unit gain cycle or tree, the result follows. If not, we can consider the following two cases.

\textbf{Case 1.} $G^{\varphi}$ contains a pendant vertex, say $x$.
Let $y$ be the unique neighbour vertex of $x$ in $G^{\varphi}$ and $$H^{\varphi}=G^{\varphi}-x-y.$$ Let $H_{1}^{\varphi},H_{2}^{\varphi},\cdots,H_{t}^{\varphi}$ be all connected components of $H^{\varphi}$. By Lemmas \ref{le:3.4}(c)
and \ref{le:3.5}, $y$ does not lie on any cycle of $G^{\varphi}$ and $H^{\varphi}$ is also $p$-lower optimal. Then
by Lemma \ref{le:3.3}(a), $H_{i}^{\varphi}$ is $p$-lower optimal for each $i\in \{1,2,\ldots,t\}$.

By induction hypothesis, we have $m(T_{H_{i}})=m(H_{i}-O(H_{i}))$ for each $i\in \{1,2,\ldots,t\}$.
Then by Lemma \ref{le:2.3}, we have that
\begin{align*}
&m(T_{G})=m(T_{H})+1\\
 &~~~~~~~~~=\sum_{i=1}^{t}m(T_{H_{i}})+1\\
 &~~~~~~~~~=\sum_{i=1}^{t}m(H_{i}-O(H_{i}))+1\\
 &~~~~~~~~~=m(H-O(H))+1\\
 &~~~~~~~~~=m(G-O(G)).
\end{align*}
Assertion (c) holds in this case.

\textbf{Case 2.} $G^{\varphi}$ has a pendant complex unit gain cycle, say $C_{q_{1}}^{\varphi}$. Let $x$ be the unique vertex with degree $3$ in $C_{q_{1}}^{\varphi}$. Let $C_{q_{1}}^{\varphi},C_{q_{2}}^{\varphi},\cdots,C_{q_{t}}^{\varphi}$ be all cycles of $G^{\varphi}$ and
$K^{\varphi}=G^{\varphi}-C_{q_{1}}^{\varphi},H^{\varphi}=K^{\varphi}+x$. By Lemma \ref{le:3.7}(f), one has that $H^{\varphi}$ is $p$-lower optimal. Because $|V(H^{\varphi})|< |V(G^{\varphi})|$, by induction hypothesis, we have
$$m(T_{H})=m(H-O(H)).$$
Note that $T_{G}\cong T_{H}$. So by Lemmas \ref{le:3.7}(d) and \ref{le:3.8}, we have
\begin{align*}
&m(T_{G})=m(T_{H})\\
 &~~~~~~~~~=m(H-O(H))\\
 &~~~~~~~~~=m(H)-\frac{\sum_{i=2}^{t}|V(C_{q_{i}})|}{2}\\
 &~~~~~~~~~=m(K)-\frac{\sum_{i=2}^{t}|V(C_{q_{i}})|}{2}\\
 &~~~~~~~~~=m(K)+\frac{|V(C_{q_{1}})|}{2}-\frac{\sum_{i=1}^{t}|V(C_{q_{i}})|}{2}\\
 &~~~~~~~~~=m(G)-\frac{\sum_{i=1}^{t}|V(C_{q_{i}})|}{2}\\
 &~~~~~~~~~=m(G-O(G)).
\end{align*}

This completes the proof.   \quad $\square$


\noindent\begin{theorem}\label{th:3.10}
Let $G^{\varphi}$ be a connected complex unit gain graph. Then $G^{\varphi}$ is $p$-upper optimal if and only if the following
three conditions all hold:
\begin{enumerate}[(a)]
 \item Any two cycles of $G^{\varphi}$ share no common vertices;
 \item Each cycle of $G^{\varphi}$ is of Type C;
 \item $m(T_{G})=m(G-O(G))$.
\end{enumerate}
\end{theorem}
\noindent\textbf{Proof.} \textbf{Sufficiency:} We will use induction on the order of $G^{\varphi}$. If $G$ is a complex unit gain
tree or cycle of Type C, the result follows from Lemmas \ref{le:2.8} and \ref{le:2.10}. If not, since $$m(T_{G})=m(G-O(G)),$$ by Lemma \ref{le:2.13}, we have that $G^{\varphi}$ contains at least one pendant vertex, say $x$. Let $y$ be the unique neighbour vertex of $x$ in $G^{\varphi}$. By Lemma \ref{le:2.13}, $y$ lies outside
any cycle of $G^{\varphi}$. Let $$H^{\varphi}=G^{\varphi}-x-y$$ and $H_{1}^{\varphi},H_{2}^{\varphi},\cdots,H_{t}^{\varphi}$ be all connected components of $H^{\varphi}$. Since $x,y$ are both outside the cycle of $G^{\varphi}$, then $$T_{H}=T_{G}-x-y$$ and
$$H-O(H)=G-O(G)-x-y.$$ By condition(c) and Lemma \ref{le:2.3}, we have
\begin{align*}
&m(T_{G})=\sum_{i=1}^{t} m(T_{H_{i}})+1\\
&~~~~~~~~~=m(G-O(G))\\
&~~~~~~~~~=\sum_{i=1}^{t} m(H_{i}-O(H_{i}))+1.
\end{align*}

Therefore $$\sum_{i=1}^{t} m(T_{H_{i}})=\sum_{i=1}^{t} m(H_{i}-O(H_{i})),$$ we observe that for each $i\in \{1,2,\ldots,t\}$, $$m(T_{H_{i}})\geq m(H_{i}-O(H_{i})),$$ so we obtain $$m(T_{H_{i}})=m(H_{i}-O(H_{i}))$$ for each $i\in \{1,2,\ldots,t\}$. Therefore, $H_{i}^{\varphi}$ satisfies (a)-(c) for each $i\in \{1,2,\ldots,t\}$.

By induction hypothesis, for each $i\in\{1,2,\ldots,t\}$,
$$p(H_{i}^{\varphi})=m(H_{i})+c(H_{i}).$$
Then by Lemmas \ref{le:2.3}, \ref{le:2.9} and \ref{le:2.11},
\begin{align*}
&p(G^{\varphi})=p(H^{\varphi})+1\\
 &~~~~~~~~=\sum_{i=1}^{t} p(H_{i}^{\varphi})+1\\
 &~~~~~~~~=\sum_{i=1}^{t} [m(H_{i})+c(H_{i})]+1\\
 &~~~~~~~~=m(H)+c(H)+1\\
 &~~~~~~~~=m(G)+c(G).
\end{align*}

\textbf{Necessity:} Let $G^{\varphi}$ be a complex unit gain graph such that $$p(G^{\varphi})=m(G)+c(G).$$ The proof for $(a)$ and
$(b)$ goes parallel as in Theorem \ref{th:3.9}, thus omitted.
We prove the assertion $(c)$ by the induction on the order of $G^{\varphi}$. If $G^{\varphi}$ is a complex unit gain tree or cycle of Type C, the result follows. If not, we can consider the following two  cases.

\textbf{Case 1.} $G^{\varphi}$ contains a pendant vertex, say $x$.
Let $y$ be the unique neighbour vertex of $x$ in $G^{\varphi}$ and $H^{\varphi}=G^{\varphi}-x-y$. Let $H_{1}^{\varphi},H_{2}^{\varphi},\cdots,H_{t}^{\varphi}$ be all connected components of $H^{\varphi}$. By Lemmas \ref{le:3.4}(c)
and \ref{le:3.5}, $y$ does not lie on any cycle of $G^{\varphi}$ and $H^{\varphi}$ is also $p$-upper optimal. Then
by Lemma \ref{le:3.3}(b), $H_{i}^{\varphi}$ is $p$-upper optimal for each $i\in\{1,2,\ldots,t\}$.
By induction hypothesis, we have
$$m(T_{H_{i}})=m(H_{i}-O(H_{i}))$$ for each $i\in \{1,2,\ldots,t\}$.
Then by Lemma \ref{le:2.3},
\begin{align*}
&m(T_{G})=m(T_{H})+1\\
&~~~~~~~~~=\sum_{i=1}^{t} m(T_{H_{i}})+1\\
&~~~~~~~~~=\sum_{i=1}^{t} m(H_{i}-O(H_{i}))+1\\
&~~~~~~~~~=m(H-O(H))+1\\
&~~~~~~~~~=m(G-O(G)).
\end{align*}

Assertion (c) holds in this case.

\textbf{Case 2.} $G^{\varphi}$ has a pendant complex unit gain cycle, say $C_{p_{1}}^{\varphi}$. Let $x$ be the unique vertex with degree $3$ in $C_{p_{1}}^{\varphi}$. Let $C_{p_{1}}^{\varphi},C_{p_{2}}^{\varphi},\cdots,C_{p_{t}}^{\varphi}$ be all cycles of $G^{\varphi}$ and $H^{\varphi}=G^{\varphi}-C_{p_{1}}^{\varphi},K^{\varphi}=H^{\varphi}+x$, where $H^{\varphi}$ is a connected component of $G^{\varphi}-x$. By Lemma \ref{le:3.4}(b), one has that
$$p(G^{\varphi}-x)=m(G-x)+c(G-x)$$
and
$$m(G-x)=m(G).$$ Then by Lemma \ref{le:3.3}(b), $H^{\varphi}$ is $p$-upper optimal. Since $C_{p_{1}}$ is odd cycle and
$m(G-x)=m(G)$, we have that $$m(G)=m(C_{p_{1}})+m(H).$$ Because $|V(H^{\varphi})|< |V(G^{\varphi})|$, by induction hypothesis, we have
$$m(T_{H})=m(H-O(H)).$$

So by Lemma \ref{le:2.4}, there exists a maximum matching $M(H)$ of $H$ such that $M(H)\cap F(H)=\emptyset$. Let $M(C_{p_{1}})$
be a maximum matching of $C_{p_{1}}$. Then $$M(G)=M(H)\cup M(C_{p_{1}})$$ is a maximum matching of
$G$, which satisfies $$M(G)\cap F(G)=\emptyset.$$ Again by Lemma \ref{le:2.4}, we get $$m(T_{G})=m(G-O(G)).$$

This completes the proof.   \quad $\square$

Using the same methods as in Theorems \ref{th:3.9} and \ref{th:3.10}, we can obtain the following two theorems.
\noindent\begin{theorem}\label{th:3.11}
Let $G^{\varphi}$ be a connected complex unit gain graph. Then $G^{\varphi}$ is $n$-lower optimal if and only if the following
three conditions all hold:
\begin{enumerate}[(a)]
\item Any two cycles of $G^{\varphi}$ share no common vertices;
\item Each cycle of $G^{\varphi}$ is of Type A;
\item $m(T_{G})= m(G-O(G))$.
\end{enumerate}
\end{theorem}

\noindent\begin{theorem}\label{th:3.12}
Let $G^{\varphi}$ be a connected complex unit gain graph. Then $G^{\varphi}$ is $n$-upper optimal if and only if the following
three conditions all hold:
\begin{enumerate}[(a)]
\item Any two cycles of $G^{\varphi}$ share no common vertices;
\item Each cycle of $G^{\varphi}$ is of Type D;
\item $m(T_{G})= m(G-O(G))$.
\end{enumerate}
\end{theorem}
\noindent\begin{remark}
 For a  complex unit gain graph $G^{\varphi}$, if $\varphi(\overrightarrow{E})\subset\{1\}$, then $G^{\varphi}$ is the underlying  graph $G$. If $\varphi(\overrightarrow{E})\subset\{1,-1\}$, then $G^{\varphi}$ is the signed graph $\Gamma$. If $\varphi(\overrightarrow{E})\subset\{1,i,-i\}$, then $G^{\varphi}$ is the mixed graph $D_{G}$. Combing with above, we know that the  results of complex unit gain graphs are also applies to simple graphs, signed graphs and mixed graphs.
\end{remark}

\end{document}